Show solution






The spring oscillates to the other end. The amplitude is 3.5m, k = 1.2, and angular frequency is = 4.3 . Solve the wave speed.

1
Expert's answer
2021-11-22T15:13:50-0500

wave number: "k=1.2"

angular frequency: "\omega=4.3"


Wavelength:




"\lambda=2\pi/k=5.2\text{ m}."

Frequency:




"f=\omega/2\pi=4.3/2\pi=0.68\text{ Hz}."



wave speed = frequency x wavelength

"=0.68Hz \u00d7 5.2m = 3.536m\/s"